# Geometric QCD I: The Hodge-Dual Surface and Quark Confinement

### v6

@@ 2 Self-dual area derivative and loop equation
The physical hypothesis of this paper is that the actual gluonic vacuum of QCD is described by this dressed solution, choosing the perturbative vacuum $W_{\text{pert}}$ as the seed. While the symmetry is exact, the separation into a perturbative factor and a purely geometric area law is physically most distinct in the large-loop limit, where the area term dominates. For small loops, the interplay between the perturbative series (and its renormalons) and the non-perturbative parameter $\kappa$ becomes complex, as discussed in Remark 5.3 regarding the Borel summation.
@@ 5 Confinement theorem
We must also specify the definition of the perturbative factor $W_{\text{pert}}[C]$ in the Wilson loop. It is uniquely defined as an asymptotic expansion generated by iterating the MM equations [12, 14], starting from $W=1$. However, the summation of this asymptotic series is not unique beyond the perturbation expansion, which proceeds in powers of the running coupling constant $\lambda_{\text{eff}}[C]\sim 1/\log(\Lambda_{\text{QCD}}|C|)$. Consequently, non-perturbative terms, such as our dressing factor, appear at double-exponential order:
$\exp\left(-\exp\left(\frac{\text{const}}{\lambda_{\text{eff}}[C]}\right)\right)$
The ambiguity associated with adding the dressing factor with an arbitrary parameter $\kappa$ corresponds to this inherent ambiguity in perturbative QCD. To be specific, we may define $W_{\text{pert}}[C]$ as the Borel sum of the gluon diagrams in the perturbative expansion. The ambiguity of this Borel sum matches the freedom in the choice of $\kappa$. We must acknowledge that our theory of confinement remains incomplete until a unique definition of $W_{\text{pert}}[C]$ is established, along with the precise relationship between $\kappa$ and the QCD mass scale $\Lambda_{\text{QCD}}$. Such a microscopic definition, leading to explicit solution for the Momentum loop $W[P]$ will be presented in [16].
@@
Subject to these caveats, we have established a precise correspondence between QCD and the area law. As demonstrated in the following section, this framework yields the standard rising Regge trajectories while avoiding the typical inconsistencies associated with string theory in four dimensions. Our approach relies solely on minimal surfaces—a classical and well-defined subject in modern geometry. While the mathematical theory of minimal surfaces as solutions to the Plateau problem is extensive, it generally does not provide explicit analytic solutions for an arbitrary loop $C$.
Fortunately, the main application of our minimal surface – the rising meson spectrum in QCD– does not require such a general solution. The relevant loops in this case, as we shall see in the [16] of this series, simplify, allowing for the analytic solution for the minimal surface and directly leading to linear Regge trajectories.
@@ 6 String Scale $\kappa$ and Gluon Condensate
The fundamental non-perturbative property of the QCD vacuum is the gluon condensate [21]. We relate the string tension $\kappa$ to this condensate by considering a circular loop $C$ of small radius $r$. According to our solution, $S[C]=2\sqrt{2}\pi r^{2}$. We compute the second area derivative at two opposite points $x_{1}=C(0)$ and $x_{2}=C(\pi)$ on the circle.
At small distances $|x_{12}|=2r$, the Operator Product Expansion (OPE) in asymptotically free QCD is dominated by the leading conformal term in $W_{pert}[C]$:
$\frac{\delta^{2}W_{pert}[C]}{\delta\sigma_{\mu\nu}(1)\delta\sigma_{\lambda\rho}(2)}$ $\propto\langle\mathrm{Tr}F_{\mu\nu}(x_{1})U_{0,\pi}F_{\lambda\rho}(x_{2})U_{\pi,2\pi}\rangle$
$\to\frac{\alpha_{eff}}{4\pi}\frac{I_{\mu\lambda}I_{\nu\rho}-I_{\nu\lambda}I_{\mu\rho}}{|x_{12}|^{4}}W_{pert}[C],$ (6.1)
where $I_{\mu\lambda}=\delta_{\mu\lambda}-2\hat{x}_{\mu}\hat{x}_{\lambda}$ with $\hat{x}=x_{12}/|x_{12}|$. We project this perturbative singularity out by contracting with the tensor $\Pi_{\mu\nu\lambda\rho}=\tfrac{1}{2}(\delta_{\mu\lambda}\delta_{\nu\rho}-\delta_{\nu\lambda}\delta_{\mu\rho})$. Simple tensor algebra confirms that the leading term vanishes: $\mathrm{Tr}(\Pi I\otimes I)=0$.
The remaining term is the gluon condensate contribution.
@@
This constant contribution must be balanced by the second area derivative of our Hodge-dual factor in the ansatz $W=W_{pert}e^{-\kappa S}$. Using the previously found area derivative (3.51) for a circle , we find:
$\frac{\delta^{2}\exp\left(-\kappa S_{\chi}[C]\right)}{\delta\sigma_{\mu\nu}(0)\delta\sigma_{\lambda\rho}(\pi)}_{\mathrm{circle}}=4\kappa^{2}\eta^{3,\chi}_{\mu\nu}\eta^{3,\chi}_{\lambda\rho}\exp\left(-\kappa S_{\chi}[C]\right)$ (6.3)
@@ 7.4 Decoupling of Confinement and Correlations
    The Target Space: In AdS/CFT and models like Sakai-Sugimoto [20], the holographic dual is a curved 5-dimensional anti-de Sitter space (plus a compact manifold), representing a gravitational theory. In our framework, the "bulk" space is a flat 12-dimensional space, $\mathbb{R}^{4}\otimes\mathfrak{su}(2)$, formed by the tensor product of Euclidean spacetime and the canonical $\mathfrak{su}(2)$ subalgebra associated with the self-dual sector.
@@
    Conformal Invariance vs. Confinement: Standard AdS/CFT begins with a conformal field theory ($\mathcal{N}=4$ SYM). To describe QCD, one must break this symmetry "by hand" via deformations of the metric, such as introducing a hard wall, a soft wall, or compactifying D4-branes on a thermal circle [23]. Our approach is the inverse: we begin with the Area Law as the fundamental physical input. The theory is inherently confining and non-conformal from the outset, with the mass scale set naturally by the string tension $\sigma$ (derived from the gluon condensate in Section 6), without ad hoc geometric deformations.

### v7

@@ 2 Self-dual area derivative and loop equation
The physical hypothesis of this paper is that the actual gluonic vacuum of QCD is described by this dressed solution, choosing the perturbative vacuum $W_{\text{fluct}}$ as the seed. While the symmetry is exact, the separation into a perturbative factor and a purely geometric area law is physically most distinct in the large-loop limit, where the area term dominates. For small loops, the interplay between the perturbative series (and its renormalons) and the non-perturbative parameter $\kappa$ becomes complex, as discussed in Remark 5.3 regarding the Borel summation.
@@ 5 Confinement theorem
We must also specify the definition of the fluctuation factor $W_{fluc}[C]$ in the Wilson loop. It is uniquely defined as an asymptotic expansion generated by iterating the MM equations [3, 4], starting from $W=1$. The physical hypothesis of this paper is that the actual gluonic vacuum of QCD is described by this dressed solution, choosing the fluctuation factor $W_{fluc}$ as the seed. Crucially, while the rigid minimal surface $S[C]$ provides the exact ground-state area law (confinement), $W_{fluc}[C]$ encapsulates all dynamic transverse gluon-exchange fluctuations. Therefore, physical phenomena such as the Lüscher term ($\pi/24R$) and the spectrum of excited flux tubes (hybrid mesons) observed in Lattice QCD are entirely generated by $W_{fluc}[C]$, perfectly complementing the geometric rigidity of the background vacuum. The ambiguity associated with adding the dressing factor with an arbitrary parameter $\kappa$ corresponds to the inherent non-perturbative ambiguity (Borel renormalons) in perturbative QCD.
5.1 The Zero Mode as the Resummed Multi-Instanton Vacuum
It is crucial to emphasize the physical meaning of the confining factor $\exp\left(-\kappa S[C]\right)$ as an exact zero mode of the loop diffusion operator. The Makeenko-Migdal loop equations represent the exact quantum equations of motion for Planar QCD. However, like any equations of motion, they do not uniquely dictate the physical vacuum state of the theory. The properties of the physical vacuum must be supplied independently, appearing mathematically in the loop calculus as the choice of a specific zero-mode background.
In standard perturbation theory, one implicitly chooses the trivial perturbative vacuum by iterating the loop hierarchy starting from the seed $W_{fluc}[C]=1$. While this successfully generates massless gluon exchange diagrams, it fails to capture confinement. To describe the true physical vacuum of QCD, the equations of motion must be solved in the presence of a nonperturbative background.
In the traditional gauge-field representation, this nonperturbative vacuum is envisioned as a complex mixed state—a dense ensemble or “liquid” of multi-instantons and anti-instantons. A fundamental difficulty in this coordinate-space representation is that any specific multi-instanton configuration $A_{\mu}^{cl}(x)$ explicitly breaks macroscopic translation invariance by localizing energy at the instanton centers. To recover the true, translation-invariant physical QCD vacuum, one must perform an exceedingly complicated superposition (integration) over the infinite-dimensional moduli space of all multi-instanton sizes, positions, and color orientations. In coordinate space, this summation is analytically intractable.
This is where the profound power of the loop space representation reveals itself. The mathematical condition for our loop-space zero mode is the self-duality (or anti-self-duality) of the area derivative, which is the exact geometric analogue of the $F_{\mu\nu}=\pm*F_{\mu\nu}$ condition for instantons. However, because our minimal area functional $S_{\chi}[C]$ depends only on the relative geometry of the boundary loop $C$ and is invariant under rigid translations, the exact translational invariance of the vacuum is preserved from the outset.
Consequently, the horrifyingly complex topological ensemble of multi-instantons in the gauge-field representation geometrically “resums” and simplifies in loop space into a single, exact, translation-invariant object: the Hodge-dual minimal surface. The Area Law is therefore not an ad hoc deformation, but rather the exact mathematical zero mode corresponding to this physical mixed-state vacuum. The string tension parameter $\kappa$, which factors out of the equations of motion entirely, acts as the macroscopic order parameter (the gluon condensate) emerging from this microscopic ensemble. In the next Section we compare the small-loop behavior of this dressed solution with the Operator Product Expansion to explicitly verify this relationship.
@@ 5.1 The Zero Mode as the Resummed Multi-Instanton Vacuum
Subject to these caveats, we have established a precise correspondence between QCD and the area law. As demonstrated in the next paper [16], this framework yields the standard rising Regge trajectories while avoiding the typical inconsistencies associated with string theory in four dimensions. Our approach relies solely on minimal surfaces—a classical and well-defined subject in modern geometry. While the mathematical theory of minimal surfaces as solutions to the Plateau problem is extensive, it generally does not provide explicit analytic solutions for an arbitrary loop $C$. Fortunately, the main application of our minimal surface – the rising meson spectrum in QCD– does not require such a general solution. The relevant loops in this case, as we shall see in the [16] of this series, simplify, allowing for the analytic solution for the minimal surface and directly leading to linear Regge trajectories.
@@ 6 String Scale $\kappa$ and Gluon Condensate
The fundamental non-perturbative property of the QCD vacuum is the gluon condensate [21]. We relate the string tension $\kappa$ to this condensate by considering a circular loop $C$ of small radius $r$.
According to our solution, for a circular loop of radius $r$ one has $S[C]=2\sqrt{2}\,\pi r^{2}$, and we evaluate the second area derivative at the two opposite points $x_{1}=C(0)$ and $x_{2}=C(\pi)$, with $|x_{12}|=2r$. In asymptotically free QCD the OPE gives, at short distances,
$\frac{\delta^{2}W_{\mathrm{pert}}[C]}{\delta\sigma_{\mu\nu}(1)\,\delta\sigma_{\lambda\rho}(2)}\propto\big\langle\mathrm{Tr}\,F_{\mu\nu}(x_{1})\,U_{0,\pi}\,F_{\lambda\rho}(x_{2})\,U_{\pi,2\pi}\big\rangle$
$\to\frac{\alpha_{\mathrm{eff}}}{4\pi}\,\frac{I_{\mu\lambda}I_{\nu\rho}-I_{\nu\lambda}I_{\mu\rho}}{|x_{12}|^{4}}\,W_{\mathrm{pert}}[C]\;+\;\cdots,$ (6.1)
where $I_{\mu\lambda}=\delta_{\mu\lambda}-2\hat{x}_{\mu}\hat{x}_{\lambda}$ and the ellipsis denotes terms less singular as $|x_{12}|\to 0$ (including local condensates). We only track the tensor structure and the first nonvanishing local operator after projection; overall Wilson-coefficient conventions match SVZ [21]. To isolate the nonperturbative dimension–4 contribution, we contract the antisymmetric index pairs with the projector $\Pi_{\mu\nu\lambda\rho}=\frac{1}{2}(\delta_{\mu\lambda}\delta_{\nu\rho}-\delta_{\nu\lambda}\delta_{\mu\rho})$. A direct contraction shows that the leading conformal structure is annihilated:
$\Pi_{\mu\nu\lambda\rho}\big(I_{\mu\lambda}I_{\nu\rho}-I_{\nu\lambda}I_{\mu\rho}\big)=0.$
Therefore the projected OPE starts at dimension 4, and the leading surviving term is the gluon condensate,
$-\frac{g^{2}}{2N_{c}}\langle(G^{a}_{\mu\nu})^{2}\rangle\,W_{\mathrm{pert}}[0].$ (6.2)
To verify this projection, define the scalar contraction
$S(d)\;:=\;\Pi_{\mu\nu\lambda\rho}\Big(I_{\mu\lambda}I_{\nu\rho}-I_{\nu\lambda}I_{\mu\rho}\Big),$
$\Pi_{\mu\nu\lambda\rho}=\tfrac{1}{2}(\delta_{\mu\lambda}\delta_{\nu\rho}-\delta_{\nu\lambda}\delta_{\mu\rho}),$
$I_{\mu\lambda}=\delta_{\mu\lambda}-2\hat{x}_{\mu}\hat{x}_{\lambda},$ (6.3)
in Euclidean dimension $d$ (so $\delta_{\mu\mu}=d$ and $\hat{x}^{2}=1$). Expanding the contraction gives
$S(d)$ $=\frac{1}{2}(\delta_{\mu\lambda}\delta_{\nu\rho}-\delta_{\nu\lambda}\delta_{\mu\rho})\big(I_{\mu\lambda}I_{\nu\rho}-I_{\nu\lambda}I_{\mu\rho}\big)$ (6.4)
$=\frac{1}{2}\left(\underbrace{\delta_{\mu\lambda}\delta_{\nu\rho}I_{\mu\lambda}I_{\nu\rho}}_{=(\mathrm{tr}\,I)^{2}}-\underbrace{\delta_{\mu\lambda}\delta_{\nu\rho}I_{\nu\lambda}I_{\mu\rho}}_{=\mathrm{tr}(I^{2})}\right)$
$+\frac{1}{2}\left(-\underbrace{\delta_{\nu\lambda}\delta_{\mu\rho}I_{\mu\lambda}I_{\nu\rho}}_{=\mathrm{tr}(I^{2})}+\underbrace{\delta_{\nu\lambda}\delta_{\mu\rho}I_{\nu\lambda}I_{\mu\rho}}_{=(\mathrm{tr}\,I)^{2}}\right)$ (6.5)
$=\mathrm{tr}\,I)^{2}-\mathrm{tr}(I^{2}).$ (6.6)
Now $\mathrm{tr}\,I=I_{\mu\mu}=d-2\hat{x}_{\mu}\hat{x}_{\mu}=d-2$, and, writing $I=\mathbf{1}-2\hat{x}\hat{x}^{T}$,
$I^{2}=(\mathbf{1}-2\hat{x}\hat{x}^{T})^{2}=\mathbf{1}-4\hat{x}\hat{x}^{T}+4\hat{x}(\hat{x}^{T}\hat{x})\hat{x}^{T}=\mathbf{1}$
$\Rightarrow\qquad\mathrm{tr}(I^{2})=d.$ (6.7)
Therefore
$S(d)=(d-2)^{2}-d=(d-1)(d-4),$
and in particular $S(4)=0$, i.e. the leading conformal tensor structure in (6) is annihilated by the $\Pi_{\mu\nu\lambda\rho}$ projection specifically in four dimensions.
The remaining term in OPE is the gluon condensate contribution.
Remark 6.1.
This projection removes the universal leading short-distance conformal tensor structure $\propto|x_{12}|^{-4}$ in (6). Subleading perturbative contributions and higher-dimensional operators are not claimed to vanish; they are less singular as $|x_{12}|\to 0$ and are suppressed by additional powers of the loop size in the small-loop expansion.
@@
This constant contribution from the gluon condensate must be balanced by the second area derivative of our Hodge-dual factor in the ansatz $W=W_{pert}e^{-\kappa S}$. Using the previously found area derivative (3.51) for a circle , we find:
@@
Remark 6.2.
All quantities in the OPE are understood at a renormalization scale $\mu\sim 1/r$. Accordingly, the Wilson coefficients, $\alpha_{s}(\mu)$, the condensate $\langle(G_{\mu\nu}^{a})^{2}\rangle_{\mu}$, and the inferred parameter $\kappa(\mu)$ are scheme- and scale-dependent; this is consistent with the intrinsic nonperturbative ambiguity in the definition of $W_{\mathrm{pert}}[C]$ discussed in previous remarks.
Remark 6.3.
In our conventions the area-law exponent is $\kappa S[C]$. Using (5.1), for a planar loop with physical Plateau area $A_{\mathrm{Pl}}[C]$ one has $S[C]=2\sqrt{2}\,A_{\mathrm{Pl}}[C]$, so the effective string tension multiplying $A_{\mathrm{Pl}}$ is
$\sigma\;=\;2\sqrt{2}\,\kappa\,.$
@@ 7.4 Decoupling of Confinement and Correlations
    The Target Space: In AdS/CFT and models like Sakai-Sugimoto [20], the holographic dual is a curved 5-dimensional anti-de Sitter space (plus a compact manifold), representing a dynamical gravitational theory. In our framework, the "bulk" space is a flat 12-dimensional space, $\mathbb{R}^{4}\otimes\mathfrak{su}(2)$, formed by the tensor product of Euclidean spacetime and the canonical $\mathfrak{su}(2)$ subalgebra associated with the self-dual sector. Because these extra dimensions represent an ’unfolded’ internal color space rather than physical spacetime, there are no propagating bulk gravitons. Consequently, pure Yang-Mills strictly avoids backreaction on this flat target space.
@@
    Conformal Invariance vs. Confinement: Standard AdS/CFT begins with a conformal field theory ($\mathcal{N}=4$ SYM). To describe QCD, one must break this symmetry "by hand" via deformations of the metric, such as introducing a hard wall, a soft wall, or compactifying D4-branes on a thermal circle [23]. Our approach is the inverse: we begin with the QCD loop equation and observe a confining factor which is left undefined by this loop equation, but physically corresponds to the unknown QCD vacuum (with some gluon condensate in it). The Area Law then follows as a mathematical zero mode of the loop diffusion operator, corresponding to the physical QCD vacuum. The theory is inherently confining and non-conformal from the outset, with the mass scale set naturally by the string tension $\sigma$ (derived from the gluon condensate in Section 6), without ad hoc geometric deformations.